# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-A Auxiliary lemmas
$\textstyle\quad+2L_{g}^{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}\|^{2}],$ (88)
@@
$\textstyle\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{t-1}\|^{2}]\leq\frac{2C_{x}}{t^{2\varsigma_{x}}}+\frac{2c_{u1}C_{z}(\lambda_{x}^{0})^{2}}{(\lambda_{z}^{0})^{2}t^{\beta_{z}+2v_{x}-2v_{z}}}$ (89)
@@
where the positive constant $d_{6}$ is given by $d_{6}=2C_{x}+2c_{u1}(C_{z}+mr\sigma_{z}^{2})(\lambda_{x}^{0})^{2}(\lambda_{z}^{0})^{-2}+2mc_{u2}(\lambda_{x}^{0})^{2}+n\sigma_{x}^{2}$.
@@
By using an argument similar to the derivation of (61), except that $\mathbb{E}[|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}|^{2}]$ in (74) is replaced with $\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]$, the right hand side of (87) satisfies
$\textstyle\mathbb{E}[\langle\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}),$ (91)
@@
$\textstyle\leq\frac{\kappa(\sigma_{g,1}^{2}+L_{g}^{2})}{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{8}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{9}}{(t+1)^{\alpha_{2}}},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))$
@@
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{1}{\kappa}\mathbb{E}[\|\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\|^{2}]$
×$\textstyle\quad\times m\mathbb{E}[\|\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{y}\bar{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))\|^{2}]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m(\sigma_{g,1}^{2}+L_{g}^{2})\sigma^{2}_{f,1}}{\kappa(t+1)},$ (92)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\left(\boldsymbol{1}_{m}\otimes\frac{1}{m}\sum_{i=1}^{m}\tilde{z}_{i}^{t}-\boldsymbol{\tilde{z}}^{t}\right)\rangle]$
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\tilde{W}\!\otimes\!I_{r})^{t}\Xi_{z}^{1}\rangle]$
$\textstyle+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\zeta}^{p}\rangle]$
$\textstyle+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\!\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle].$ (93)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\tilde{W}\otimes I_{r})^{t}\Xi_{z}^{1}\rangle]$ (94)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{4d_{10}}{\kappa(t+1)^{2-2v_{z}}},$
@@
Using an argument similar to the derivation of (79), except that $\mathbb{E}[|\boldsymbol{\bar{x}}^{t}-\boldsymbol{\bar{x}}^{t-1}|^{2}]$ in (74) is replaced with $\mathbb{E}[|\Xi_{x}^{t}-\Xi_{x}^{t-1}|^{2}]$, the second term on the right hand side of (93) satisfies
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(x^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\zeta}^{p}\rangle]$ (95)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle]$
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}\big{[}\big{\langle}\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}$
×$\textstyle\quad\times\lambda_{z}^{p}\left(\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p})-\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p-1})\right)\big{\rangle}\big{]},$ (96)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$ (97)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
By substituting (42) and (89) into (98), we have
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$ (100)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{14}}{\kappa(t+1)^{\beta_{1}}},$
@@
$\textstyle\mathbb{E}\left[\left\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}\right.\right.$
×$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes\left(\frac{1}{\lambda_{y}^{p}}(\bar{y}^{p+1}-\bar{y}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\bar{y}^{p}\!-\!\bar{y}^{p-1})\right)\right\rangle\right]$
$=\textstyle\mathbb{E}\left[\left\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}\right.\right.$
×$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes(g^{p}(\boldsymbol{x}^{p})-g^{p-1}(\boldsymbol{x}^{p-1}))\right\rangle\right]$
$\textstyle\quad+\mathbb{E}\left[\left\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}\right.\right.$
×$\left.\left.\textstyle\quad\times\left(\boldsymbol{1}_{m}\otimes\left(\frac{\bar{\chi}_{w}^{p}}{\lambda_{y}^{p}}-\frac{\bar{\chi}_{w}^{p-1}}{\lambda_{y}^{p-1}}\right)\right)\right\rangle\right].$ (103)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
×$\textstyle\quad\times\boldsymbol{1}_{m}\otimes(g^{p}(\boldsymbol{x}^{p})-g^{p-1}(\boldsymbol{x}^{p-1}))\rangle]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}}{\kappa(t+1)^{\beta_{1}}},$ (104)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}\lambda_{z}^{p}\Upsilon_{\chi}^{p}\rangle]$
@@
Substituting (104) and (105) into (103) yields
$\textstyle\mathbb{E}\left[\left\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}\right.\right.$
×$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes\left(\frac{1}{\lambda_{y}^{p}}(\bar{y}^{p+1}-\bar{y}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\bar{y}^{p}\!-\!\bar{y}^{p-1})\right)\right\rangle\right]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}}{(t+1)^{\alpha_{2}}}.$ (106)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
×$\textstyle\quad\times(\frac{1}{\lambda_{y}^{p}}(\boldsymbol{\hat{y}}^{p+1}-\boldsymbol{\hat{y}}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\boldsymbol{\hat{y}}^{p}-\boldsymbol{\hat{y}}^{p-1}))\rangle]$
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
×$\textstyle\quad\times\lambda_{z}^{p}\left((\tilde{W}\otimes I_{r})^{p-1}\big{(}\frac{1}{\lambda_{y}^{2}}(\boldsymbol{\hat{y}}^{2}-\boldsymbol{\hat{y}}^{1})-\frac{1}{\lambda_{y}^{1}}(\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0})\big{)}\right)$
$\textstyle\quad+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
×$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}(\Theta_{\chi}^{k}+\Theta_{g}^{k})\rangle].$ (110)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
×$\textstyle\quad\times\lambda_{z}^{p}\left((\tilde{W}\otimes I_{r})^{p-1}\big{(}\frac{1}{\lambda_{y}^{2}}(\boldsymbol{\hat{y}}^{2}-\boldsymbol{\hat{y}}^{1})-\frac{1}{\lambda_{y}^{1}}(\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0})\big{)}\right)$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{21}}{\kappa(t+1)^{2}},$ (111)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$ (112)
×$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{\chi}^{k}\rangle]$

### v3

@@ VII-A Auxiliary lemmas
$\textstyle\quad+2L_{g}^{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}\|^{2}],$ (82)
@@
$\textstyle\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{t-1}\|^{2}]\leq\frac{2\bar{C}_{x}}{t^{2\varsigma_{x}}}+\frac{2c_{u1}\bar{C}_{z}(\lambda_{x}^{0})^{2}}{(\lambda_{z}^{0})^{2}t^{\beta_{z}+2v_{x}-2v_{z}}}$ (83)
@@
where the positive constant $d_{6}$ is given by $d_{6}=2\bar{C}_{x}+2c_{u1}(\bar{C}_{z}+mr\sigma_{z}^{2})(\lambda_{x}^{0})^{2}(\lambda_{z}^{0})^{-2}+2mc_{u2}(\lambda_{x}^{0})^{2}+n\sigma_{x}^{2}$.
@@
By using an argument similar to the derivation of (55), except that $\mathbb{E}[|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}|^{2}]$ in (68) is replaced with $\mathbb{E}[\|\Xi_{x}^{t}-\Xi_{x}^{t-1}\|^{2}]$, the right hand side of (81) satisfies
$\textstyle\mathbb{E}[\langle\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})^{\top}(\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}),$ (85)
@@
$\textstyle\leq\frac{\kappa(\sigma_{g,1}^{2}+L_{g}^{2})}{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{8}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{9}}{(t+1)^{\alpha_{2}}},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t})))$
@@
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{1}{\kappa}\mathbb{E}[\|\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\|^{2}]$
$\textstyle\quad\times m\mathbb{E}[\|\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{y}\bar{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))\|^{2}]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m(\sigma_{g,1}^{2}+L_{g}^{2})\sigma^{2}_{f,1}}{\kappa(t+1)},$ (86)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\left(\boldsymbol{1}_{m}\otimes\frac{1}{m}\sum_{i=1}^{m}\tilde{z}_{i}^{t}-\boldsymbol{\tilde{z}}^{t}\right)\rangle]$
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\tilde{W}\!\otimes\!I_{r})^{t}\Xi_{z}^{1}\rangle]$
$\textstyle+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{1}\Xi_{\zeta}^{p}\rangle]$
$\textstyle+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\!\sum_{p=1}^{t}(\tilde{W}\!\otimes\!I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle].$ (87)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})(\tilde{W}\otimes I_{r})^{t}\Xi_{z}^{1}\rangle]$ (88)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{4d_{10}}{\kappa(t+1)^{2-2v_{z}}},$
@@
Using an argument similar to the derivation of (73), except that $\mathbb{E}[|\mathbf{\bar{x}}^{t}-\mathbf{\bar{x}}^{t-1}|^{2}]$ in (68) is replaced with $\mathbb{E}[|\Xi_{x}^{t}-\Xi_{x}^{t-1}|^{2}]$, the second term on the right hand side of (87) satisfies
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(x^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{1}\Xi_{\zeta}^{p}\rangle]$ (89)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle]$
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$
$\textstyle\quad+\mathbb{E}\big[\big\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}$
$\textstyle\quad\times\lambda_{z}^{p}\left(\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p})-\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p-1})\right)\big\rangle\big],$ (90)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$ (91)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
By substituting (83) into (92), we have
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Upsilon_{f}^{p}\rangle]$ (94)
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{14}}{\kappa(t+1)^{\beta_{1}}},$
@@
[…]
$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes\left(\frac{1}{\lambda_{y}^{p}}(\bar{y}^{p+1}-\bar{y}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\bar{y}^{p}\!-\!\bar{y}^{p-1})\right)\right\rangle\right]$
[…]
$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes(g^{p}(\boldsymbol{x}^{p})-g^{p-1}(\boldsymbol{x}^{p-1}))\right\rangle\right]$
[…]
$\left.\left.\textstyle\quad\times\left(\boldsymbol{1}_{m}\otimes\left(\frac{\bar{\chi}_{w}^{p}}{\lambda_{y}^{p}}-\frac{\bar{\chi}_{w}^{p-1}}{\lambda_{y}^{p-1}}\right)\right)\right\rangle\right].$ (97)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
$\textstyle\quad\times\boldsymbol{1}_{m}\otimes(g^{p}(\boldsymbol{x}^{p})-g^{p-1}(\boldsymbol{x}^{p-1}))\rangle]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}}{\kappa(t+1)^{\beta_{1}}},$ (98)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}\lambda_{z}^{p}\Upsilon_{\chi}^{p}\rangle]$
@@
Substituting (98) and ($9$) into (97) yields
[…]
$\left.\left.\textstyle\quad\times\boldsymbol{1}_{m}\otimes\left(\frac{1}{\lambda_{y}^{p}}(\bar{y}^{p+1}-\bar{y}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\bar{y}^{p}\!-\!\bar{y}^{p-1})\right)\right\rangle\right]$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}}{(t+1)^{\alpha_{2}}}.$ (100)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
$\textstyle\quad\times(\frac{1}{\lambda_{y}^{p}}(\boldsymbol{\hat{y}}^{p+1}-\boldsymbol{\hat{y}}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\boldsymbol{\hat{y}}^{p}-\boldsymbol{\hat{y}}^{p-1}))\rangle]$
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
$\textstyle\quad\times\lambda_{z}^{p}\left((\tilde{W}\otimes I_{r})^{p-1}\big(\frac{1}{\lambda_{y}^{2}}(\boldsymbol{\hat{y}}^{2}-\boldsymbol{\hat{y}}^{1})-\frac{1}{\lambda_{y}^{1}}(\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0})\big)\right)$
$\textstyle\quad+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}(\Theta_{\chi}^{k}+\Theta_{g}^{k})\rangle].$ (104)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
$\textstyle\quad\times\lambda_{z}^{p}\left((\tilde{W}\otimes I_{r})^{p-1}\big(\frac{1}{\lambda_{y}^{2}}(\boldsymbol{\hat{y}}^{2}-\boldsymbol{\hat{y}}^{1})-\frac{1}{\lambda_{y}^{1}}(\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0})\big)\right)$
$\textstyle\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{21}}{\kappa(t+1)^{2}},$ (105)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$ (106)
$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{\chi}^{k}\rangle]$